Calculate the extent of reaction at which gelation occurs for the following mixtures of glycerol/terephthalic acid/trimesic acid (using the Carothers approach); ‘unnormalized’ molar ratios are provided below: (f) 1.0/1.0/0 (g) 1.0/0/1.0 (h) 1.0/0.5/0.5 (i) 1.5/0.5/0.5 (j) 1.5/0.99/0.002
Answer #1

According to Carothers approach, the extent or steaction at which gelation occurs is given by the formula — f - I Ni where ,$ = 2.5 f= 1.0x3+1.0 x2 +0x3 1.0+ 1.0 +0 :: P=2.5 10.8) 0 2,5 = 3 (g) +- 1.0x3 + 0x2 + 1.0x3 _ z 1.0 + 0 +1.0 :. PŞ-.67] cht